OpenStudy (czarluc):

sec(x)+tan(x)=0 why does it have no solution?

OpenStudy (anonymous):

because, sec(x) + tan(x) = 1 1/cos(x) + sin(x)/cos(x) = 1 (1 + sin(x)) / cos(x) = 1 1 + sin(x) = cos(x) 1 = cos(x) - sin(x) x = 0 radians, 2π, 4π ... 2nπ radians

OpenStudy (czarluc):

but the equation is equal to zero but your's is equal to one?

OpenStudy (anonymous):



OpenStudy (anonymous):

can u understand this?